0 0
Add a comment Improve this question Transcribed image text
Answer #1

Answer

A firm hires both inputs at equal marginal product per dollar

\frac{MPL}{wage}=\frac{MPK}{r}

MPL=\frac{\partial Q}{\partial L}=\frac{\partial \left ( 400L-20L^2+600K-10K^2 \right )}{\partial L}=400-40L\\ MPL=\frac{\partial Q}{\partial K}=\frac{\partial \left ( 400L-20L^2+600K-10K^2 \right )}{\partial K}=600-20K\\ \frac{400-40L}{20}=20-2L=\frac{600-20K}{30}=20-\frac{2}{3}K\\ 20-2L=20-\frac{2}{3}K\\ L=\frac{1}{3}K\\ from\ cost\\ wL+rK=TC\\ 20L+30K=550\\ 20*\frac{1}{3}K+30K=550\\ \frac{110}{3}K=550\\ K=15

the frim should hire 15 units of capital.